OpenStudy (anonymous):

IM SERIOUSLY ABOUT TO START CRYING! SOMEONE PLEASE HELP ME IM SO FRUSTRATED!!!!

OpenStudy (anonymous):

Log 5^(2x-3)=18

OpenStudy (anonymous):

@invinceble @undeadknight26 PLEASE HELP ME!

OpenStudy (phi):

is this \[ \log_{10}\left( 5^{2x-3}\right) = 18 \] or something else ?

OpenStudy (anonymous):

yes, that is it, i tried to make it like that but it would not work.

OpenStudy (phi):

So it is log base 10? do you know this "rule" \[ \log(a^b) = b \ \log(a) \]?

OpenStudy (anonymous):

no i dont know anything about logs.. Im doing a dba, and i was supposed to get out of school over 30 mins ago... i have 2 questions left..

OpenStudy (phi):

what is a dba?

OpenStudy (anonymous):

Discussion based assessment.

OpenStudy (anonymous):

messaging my teach about logs.

OpenStudy (phi):

logs are confusing but you can survive if you memorize some rules start with the one just posted \[ \log(a^b) = b \ \log(a) \] do you see how to use that rule to re-write your problem?

OpenStudy (anonymous):

no.

OpenStudy (phi):

pattern match \[ \log_{10}\left( 5^{2x-3}\right) \\ \log_{10}(a^b) \]

OpenStudy (phi):

the "a" in the rule matches with what ? and the "b" matches with what (in the problem) ?

OpenStudy (anonymous):

a=5 b=^2x-3

OpenStudy (phi):

b is just the (2x-3) the ^ is just how to show the 2x-3 is an exponent now use the rule \[ \log(a^b) = b \ \log(a) \] what do we get ?

OpenStudy (anonymous):

i have no idea...

OpenStudy (phi):

the rule says if you have a^b inside a log, you can "move the b" outside the law and multiply.

OpenStudy (phi):

Use a=5 b=2x-3 and write b log(a) using those expressions for a and b

OpenStudy (anonymous):

so... UGH im still confused. -.-

OpenStudy (phi):

You are learning how to use the rule \[ \log(a^b) = b \log(a) \] (learning why it works is possible, but I am guessing you don't want to learn that) to use the rule you match your problem with the left side you did that. you found a=5 and b= (2x-3) now use the right side. write b log(a) but use 5 for a, and 2x-3 for b can you try ?

OpenStudy (anonymous):

I dont have time...

OpenStudy (anonymous):

i have to go now.. >.< my teacher is calling me... :/ Thank you anyways.

OpenStudy (phi):

ok, next time ask before the dba